OpenStudy (anonymous):

Select the fraction that is equivalent to 0.0368.

OpenStudy (radar):

\[368\over10000\] Can be reduced if needed

OpenStudy (turingtest):

368/10000=184/5000=97/2500

OpenStudy (radar):

\[23\over1250\]

OpenStudy (turingtest):

No radar 23/625

OpenStudy (radar):

Thats better. both numerator and denominator can be divided by 16, somewhere I got lost, but so did you in that97/2500. Tricky hum?